OpenStudy (anonymous):

how many solutions does the system of equations have? 2x+3y=6 y=-2/3x-1

OpenStudy (anonymous):

2x+3y=6 --> y = (-2/3)x + 2 y=-2/3x-1 ---------------- The 2 lines are parallel if you graph them. No intersection, no solution. 2x+3y=6 --> y = (-2/3)x + 2 y=-2/3x-1 ---------------- The 2 lines are parallel if you graph them. No intersection, no solution.